# FreeQ-Graph: Free-form Querying with Semantic Consistent Scene Graph for 3D Scene Understanding

### v1

@@ II-1 Open-Vocabulary 3D Scene Understanding
Natural language querying in complex 3D scenes demands deep understanding of free-form semantics and relationships. Many prior works [11, 10, 17, 18, 19, 20, 21] rely on joint training with large-scale pretrained data to align 3D scenes and query embeddings, but their dependence on predefined vocabularies limits true free-form querying. Recent advanced methods [14, 13, 6, 22, 23] leverage large language models (LLMs) for flexible semantic reasoning, yet they overly depend on LLMs and LVLMs to generate semantic labels for 3D features without sufficient 3D scene awareness. This often results in inconsistent or inaccurate outputs where objects and relations misalign with actual 3D semantics. Additionally, some LLM-based methods [3, 24, 25, 19] fine-tune on task-specific datasets, improving performance on those tasks but still restricting free-form queries and requiring substantial training resources. Our work utilizes a training-free, free-form querying framework that constructs a scene graph with accurate nodes and relations, aligns 3D instances with correct semantics for scene understanding.
@@
The 3D Scene Graph (3DSG) represents scene semantics compactly, with nodes for objects and edges for their relationships [26, 27, 28]. Recent methods [29, 30] use 3DSG for 3D scene representations. These works [31, 32, 26], such as VL-SAT [31], construct scene graphs to model scene representations, but are constrained by the closed vocabularies from their training data, limiting their ability to support free-form semantic queries. More recent approaches like ConceptGraph [13] and BBQ [14] leverage LLMs to generate nodes and edges in scene graphs. However, their heavy reliance on LLM-generated outputs without incorporating 3D scene context often leads to inconsistent scene representations misaligned with actual 3D semantics. Our approach constructs a semantically consistent 3D scene graph by first obtaining complete and accurate free-form nodes, then aligning them with correct semantic labels.
@@ III-A Problem Formulation
3D semantic consistent feature generation. We extract visual descriptors (CLIP [7]) from 2D masks. Additionally, we generate 3D instance semantic labels via superpoint clustering and align nodes with point-level semantic representations (see Sec. III-C for details). The final nodes $\mathbf{v}_{i}$ consist of point cloud $\mathbf{p}_{i}$, unit-normalized feature ${{\hat{\mathbf{f}_{i}}}}$, and 3D box $\mathbf{b}_{i}$.
3D nodes caption generation. For each posed image, building on ConceptGraph [13], we generate node captions via LVLM+LLM: (1) prompt LVLM with “describe the central object” at top-$n$ clean viewpoints for initial descriptions; (2) distill coherent captions $\mathbf{c}_{i}$ via LLM refinement.
@@
3D semantic aligned features. After obtaining 3D semantic feature labels, we aim to align scene graph nodes with point-level semantic representations for 3D semantic aligned features. Unlike prior works [13, 14] that rely solely on a visual encoder for per-point representation, we employ both a vision encoder and a text encoder to ensure semantic consistency between nodes and point representations. Specifically, for each the node $\mathbf{v}_{i}$ and its assigned semantic superpoint label $Q_{i}$, we first extract its visual feature $\mathbf{f}_{i}$ using CLIP. Additionally, we encode the superpoint’s semantic label $Q_{i}$ through a encoder to obtain its semantic feature ${{\mathbf{f}_{Q_{i}}}}$, and fuse it with the visual feature via meanpooling $\varphi$ to obtain the final semantically aligned representation ${\hat{\mathbf{f}_{i}}}$.
@@
where $\varphi$ is the mean pooling. The ${{\hat{\mathbf{f}_{i}}}}$ is used to align semantic labels (e.g., “desk”) to nodes in the graph. Errors in LLMs and LVLMs may misassign labels of 3D instance features, leading to incorrect results in JSON-based reasoning.
@@ IV-B Experiment Results
Besides, for Sr3D and Nr3D datasets, we evalute the top-1 using groud-truth boxes in Table II and the accuray at 0.1 and 0.25 IOU threshold for 5 categories in Table III. As shown in Table II, we validated the top-1 performance using ground-truth boxes across fully-supervised models [17, 10, 6], weakly-supervised model [55], and zero-shot models [23, 49]. Our model achieved best results across 5 different metrics, demonstrating its superior performance. As shown in Table III, our model also consistently outperformed all SOTA works [63, 13, 14] across 4 cases of Sr3D and Nr3D datasets. Compared to ConceptGraph [13], BBQ [14] and Open3DSG [32], which also utilize LLMs and graph representations for reasoning, our model shows significant advantages, validating the semantic aligned features in reasoning with free-form queries. Fig.4 shows the quantitative comparison. SeeGround [23] fails to capture object relationships like “near”, while BBQ [14] struggles with semantic label like “single sofa”, hindering accurate grounding. In contrast, our model precisely grounds objects with correct semantic labels and understands both scene-level and object-level spatial relationships.
@@ IV-B2 Complex Queries
Table V shows our 3D scene graph evaluation on 3DSSG [28], where we surpass all SOTA models [62, 13, 14, 32] in object, predicate, and relationship prediction.This demonstrates that methods leverage on pre-train datasets, such as OpenSeg [62], are unsuitable for predicting object nodes and relationships, while models over-rely on LLMs(e.g., ConceptGraph [13], BBQ [14]) fail on small or thin objects. While Open3DSG [32] can predict open-vocabulary objects, it still faces challenges in free-form relationship prediction. In contrast, our model can precisely predict free-form objects, predicates, and relationships without any training priors. As shown in Fig 6, it illustrates that our model not only creates a semantically consistent scene graph with complete nodes and correct relations but also assigns the correct semantic labels to each node.

### v2

@@ II-1 Open-Vocabulary 3D Scene Understanding
Natural language querying in complex 3D scenes demands deep understanding of free-form semantics and relationships. Many prior works [11, 10, 17, 18, 19, 20, 21, 22, 23] rely on joint training with large-scale pretrained data to align 3D scenes and query embeddings, but their dependence on predefined vocabularies limits true free-form querying. Recent advanced methods [14, 13, 6, 24, 25] leverage large language models (LLMs) for flexible semantic reasoning, yet they overly depend on LLMs and LVLMs to generate semantic labels for 3D features without sufficient 3D scene awareness [26]. This often results in inconsistent or inaccurate outputs where objects and relations misalign with actual 3D semantics. Besides, some LLM-based methods [3, 27, 28, 19] fine-tune on task-specific datasets, improving performance but still restricting free-form queries and requiring substantial training resources. Our work utilizes a training-free, free-form querying framework that constructs a scene graph with accurate nodes and relations, aligns 3D instances with correct semantics for scene understanding.
@@
The 3D Scene Graph (3DSG) represents scene semantics compactly, with nodes for objects and edges for their relationships [29, 30, 31, 32]. Recent methods [33, 34, 35] use 3DSG for 3D scene representations. These works [36, 37, 29], such as VL-SAT [36], construct scene graphs to model scene representations, but are constrained by the closed vocabularies from their training data, limiting their ability to support free-form semantic queries. More recent approaches like ConceptGraph [13] and BBQ [14] leverage LLMs to generate nodes and edges in scene graphs. However, their heavy reliance on LLM-generated outputs without incorporating 3D scene context often leads to inconsistent scene representations misaligned with actual 3D semantics. Our approach constructs a semantically consistent 3D scene graph by first obtaining complete and accurate free-form nodes, then aligning them with correct semantic labels.
@@ III-A Problem Formulation
3D semantic consistent feature generation. We extract visual descriptors (CLIP [7]) from 2D masks. Additionally, we generate 3D instance semantic labels via superpoint clustering and align nodes with point-level semantic representations (see Sec. III-C for details). The final nodes $\mathbf{v}_{i}$ consist of point cloud $\mathbf{p}_{i}$, unit-normalized feature $^{{\hat{\mathbf{f}_{i}}}}$, and 3D box $\mathbf{b}_{i}$.
[…]
@@
3D semantic aligned features. After obtaining 3D semantic feature labels, we aim to align scene graph nodes with point-level semantic representations for 3D semantic aligned features. Unlike prior works [13, 14] that rely solely on a visual encoder for per-point representation, we employ both a vision encoder and a text encoder to ensure semantic consistency between nodes and point representations. Specifically, for each the node $\mathbf{v}_{i}$ and its assigned semantic superpoint label $Q_{i}$, we first extract its visual feature $\mathbf{f}_{i}$ using CLIP. Additionally, we encode the superpoint’s semantic label $Q_{i}$ through a encoder to obtain its semantic feature ${{\mathbf{f}_{Q_{i}}}}$, and fuse it with the visual feature via meanpooling $\varphi$ to obtain the final semantically aligned representation $^{\hat{\mathbf{f}_{i}}}$.
@@
where $\varphi$ is the mean pooling. The $^{{\hat{\mathbf{f}_{i}}}}$ is used to align semantic labels (e.g., “desk”) to nodes in the graph. Errors in LLMs and LVLMs may misassign labels of 3D instance features, leading to incorrect results in JSON-based reasoning.
@@ IV-B Experiment Results
Besides, for Sr3D and Nr3D datasets, we evalute the top-1 using groud-truth boxes in Table II and the accuray at 0.1 and 0.25 IOU threshold for 5 categories in Table III. As shown in Table II, we validated the top-1 performance using ground-truth boxes across fully-supervised models [17, 10, 6], weakly-supervised model [60], and zero-shot models [25, 54]. Our model achieved best results across 5 different metrics, demonstrating its superior performance. As shown in Table III, our model also consistently outperformed all SOTA works [68, 13, 14] across 4 cases of Sr3D and Nr3D datasets. Compared to ConceptGraph [13], BBQ [14] and Open3DSG [37], which also utilize LLMs and graph representations for reasoning, our model shows significant advantages, validating the semantic aligned features in reasoning with free-form queries. Fig.4 shows the quantitative comparison. SeeGround [25] fails to capture object relationships like “near”, while BBQ [14] struggles with semantic label like “single sofa”, hindering accurate grounding. In contrast, our model precisely grounds objects with correct semantic labels and understands both scene and object-level spatial relationships.
@@ IV-B2 Complex Queries
As shown in Table V, we evaluate our 3D scene graph generation on the 3DSSG [31]. Our method outperforms most fully-supervised models and all zero-shot baselines [67, 13, 14, 37], while maintaining a training-free and highly efficient pipeline. Although our model does not surpass VL-SAT, which is trained on large-scale vision-language datasets, it achieves competitive performance without requiring any training. Moreover, in contrast to existing methods that over-rely on LLMs (e.g., ConceptGraph [13], BBQ [14]), our approach demonstrates robustness in predicting small or thin objects, where these methods typically fails. As shown in Fig 6, it illustrates that our model not only creates a semantically consistent scene graph with complete nodes and correct relations but also assigns the correct semantic labels to each node.